OpenStudy (anonymous):

Which is equal to (√2)^5?

OpenStudy (anonymous):

Options: \[2\frac{ 1 }{ 10 }\] \[2\frac{ 2 }{ 5 }\] \[2\frac{ 5 }{ 2 }\] \[2^{10}\]

OpenStudy (vishweshshrimali5):

Ok

OpenStudy (anonymous):

I think it's D, \[2^{10}\]

OpenStudy (vishweshshrimali5):

Have you given it some try ?

OpenStudy (anonymous):

Because 5*2=10 so 2 is to the 10th power

OpenStudy (vishweshshrimali5):

No, actually it is 1/2 * 5 = 5/2

OpenStudy (anonymous):

I suck at algebra (can you tell?) I really wish I got to take 2 years of geometry instead of this lol

OpenStudy (vishweshshrimali5):

Its okay

OpenStudy (anonymous):

So it'd be \[2\frac{ 5 }{ 2 }\]

OpenStudy (vishweshshrimali5):

See: by law of exponents we have: 

OpenStudy (vishweshshrimali5):

Now, here, a = 2 and m = 1/2 (because square root means power 1/2 ) and n is ?

OpenStudy (vishweshshrimali5):

Can you guess ?

OpenStudy (anonymous):

1?

OpenStudy (vishweshshrimali5):

Well I would be back in 1 hour.

OpenStudy (vishweshshrimali5):

No, n is 5 here

OpenStudy (tkhunny):

\(\sqrt{2}^{5} = \sqrt{2}^2\cdot\sqrt{2}^2\cdot \sqrt{2}\)

OpenStudy (anonymous):

Thank you @vishweshshrimali5 and @tkhunny for answering!

OpenStudy (tkhunny):

Are you sure we're not supposed to be talking about \(\left(2^{2}\right)^{5}\)?